# Phased Instruction Fine-Tuning for Large Language Models

### v1

@@ 4.2 Experiment 2: Comparison win rate of difficulty-stratified 3-stages and randomly sampled 3-stages
Permutations Self-Instruct WizardLM Koala Vicuna OASST Anthropic Base Models
1-2-3 +5.75 +10.32 +4.17 +11.25 +8.78 +12.40 7B
2-1-3 +0.40 +2.98 +0.56 +11.25 +6.65 +8.14
3-1-2 +3.57 +1.15 -3.33 +12.50 +1.06 +5.43
1-3-2 -0.79 -2.52 -0.83 +2.50 -2.39 +5.43
2-3-1 -5.16 -9.40 -13.61 -8.12 -5.32 -5.04
3-2-1 -4.76 -5.96 -9.17 -6.25 -9.04 -1.16
1-2-3 +4.37 +6.88 +10.56 +11.25 +10.11 +13.18 13B
2-1-3 +6.15 +7.11 +10.28 +5.62 +5.85 +8.53
3-1-2 +0.79 +1.38 +4.44 -6.88 +2.39 +2.33
1-3-2 +1.39 +3.90 +5.00 +1.25 +1.06 +6.98
2-3-1 -3.37 -6.65 -1.11 -7.50 -11.70 -10.08
3-2-1 -5.16 -5.28 -5.00 -11.88 -12.77 -7.36
@@
4.3 Experiment 3: Comparison win rate of uptraining on Alpaca-3-stages in all permutations
@@
To investigate the impact of varying the sequence of Alpaca-3-stages on the performance of the Llama-2 7B/13B models with clarity and precision, the following structure is proposed. This structure presents the uptraining order across the three stages, denoted as Stage 1 (1), Stage 2 (2), and Stage 3 (3), to illustrate the sequence in which the model was uptrained on different segments of the instruction data:
@@
Table 4 demonstrates the performance enhancements achieved by fine-tuning the pretrained Llama-2 7B/13B models through the Phased IFT method across 6 permutations of these stages, in contrast to the One-off IFT method applied directly on the original Alpaca data. The results distinctly reveal that the Llama-2 7B/13B models achieve maximal win-rate improvements when sequentially fine-tune in the 1-2-3 order using Phased IFT, substantially outperforming other sequential arrangements.
The Llama-2 7B model, after undergoing Phased IFT in a 1-2-3 sequence, demonstrates significant win-rate enhancements of +5.75, +10.32, +4.17, +11.25, +8.78, and +12.40 absolute points across Self-Instruct, WizardLM, Koala, Vicuna, OASST, and Anthropic benchmarks, respectively. This outperformance is substantial when compared to the model fine-tuned via the One-off IFT method on the 52K Alpaca data, as well as models fine-tuned in alternative sequences. Notably, the 1-2-3 and 2-1-3 training sequences yield markedly superior win rate than the 3-1-2 and 1-3-2 sequences, which in turn outperform the 2-3-1 and 3-2-1 sequences. The latter sequences exhibit win rate significantly lower than those achieved through One-off IFT.
Similarly, the Llama-2 13B model, fine-tuned in the 1-2-3 sequence using Phased IFT, show win-rate improvements of +4.37, +6.88, +10.56, +11.25, +10.11, and +13.18 absolute points on the same benchmarks, respectively, again surpassing the performance of models fine-tuned through other sequences or the One-off IFT method. The training sequences of 1-2-3 and 2-1-3 consistently resulted in higher win rate compared to 3-1-2 and 1-3-2, which outperform 2-3-1 and 3-2-1 sequences. The win rate for the 2-3-1 and 3-2-1 sequences are notably inferior to those achieved with One-off IFT.
These findings compellingly support the hypothesis that uptraining in order of increasing difficulty presents significant positive gains, whereas reverse order training yields negligible benefits. This underpins our progressive alignment hypothesis, positing that the process of aligning pretrained models to human intent through supervised instruction-based fine-tuning is inherently gradual. Beginning with basic and straightforward instructions before progressively tackling more challenging ones ensures a systematic alignment with end tasks. This approach not only validates the utility of constructing diverse and difficulty-tiered instruction data but also highlights the strategic importance of adopting a progressively challenging fine-tuning regimen to enhance model performance and alignment with human intent in task completion.
To summarize, Experiment 1 validates the superiority of phased instruction fine-tuning (Phased IFT) over one-off IFT, evidenced by a notable enhancement in win rates. Experiment 2 underscores the significance of stratifying instruction dataset by difficulty score, yielding a pronounced improvement in win rate against random sampling for uptraining fine-tuning, as depicted in Row #3 of Table 3, with the former method achieving a significant uplift in win rate. Experiment 3 demonstrates that sequential uptraining fine-tuning by ascending difficulty represents the optimal phased training approach, concurrently corroborating the progressive alignment hypothesis. Collectively, these experiments show the efficacy of the progressive alignment hypothesis and Phased IFT.
@@ 5 Conclusion
The paper introduces the progressive alignment hypothesis, suggesting that the transition of a pretrained model’s next-word prediction capability into an ability to follow human intentions is a sequential learning process. This process entails progressively aligning with human intent, advancing from simpler to more complex tasks. In light of this hypothesis, we propose a phased instruction fine-tuning (Phased IFT) approach. Utilizing GPT-4 to score the difficulty of instructions, we categorize the instruction dataset into different stages according to the difficulty score density distribution. Diverging from prior methods that apply a one-off instruction fine-tuning (One-off IFT) on the whole instruction data, our approach entails staged, supervised uptraining, addressing each stage in ascending order of difficulty. Through extensive experiments, we have validated the progressive alignment hypothesis and the efficacy of Phased IFT, demonstrating its utility in developing diverse instruction dataset and enhancing instruction fine-tuning.
@@ 6 Broader Impact
This paper discusses a fundamental issue in Large Language Models (LLMs): how to align LLMs more efficiently. The work introduces a simple and effective method for alignment, which involves analyzing the difficulty of instruction dataset and structuring the dataset into several stages, followed by supervised fine-tuning through uptraining. We conduct three experiments, with an additional experiment detailed in Appendix A.3, to actively validate the existence of the progressive alignment hypothesis. If the progressive alignment hypothesis is a pervasive phenomenon, this work is poised to have a widespread impact on the field of LLMs.
The datasets employed in this work, namely Alpaca 52K and Alpaca-cleaned, are publicly available, and our training framework leverages open-source tools, utilizing Hugging Face’s Trainer and DeepSpeed. Furthermore, we have purchased access to OpenAI’s API to employ GPT-4 for assessing the difficulty of instruction datasets and for evaluating the quality of generated responses. We will make all our code and data openly available, including training code, GPT-4 difficulty scoring code, win-rate calculation code, and the difficulty-stratified Alpaca and Alpaca-cleaned instruction datasets.
@@ 7 Limitations
Automatic instruction dataset segmentation This work employs heuristic information from the cumulative density curve of difficulty scores, supplemented by human judgment for threshold selection. A significant challenge lies in developing algorithms capable of autonomously identifying threshold points from this curve, eliminating the reliance on human intuition and ensuring balanced data size across multi-stages sub-datasets.
Validation on larger-scale LLMs Our empirical analysis is confined to pre-trained models with capacities of 7B and 13B, lacking empirical validation on larger-scale models, such as 65B or 70B. This limitation suggests a necessity for extending validation efforts to encompass a wider range of model sizes to enhance the generalizability of our findings.
@@ Appendix A Appendix
A.1 Histogram and density curve for Alpaca, Alpaca-cleaned and LIMA
@@
We conducted a comparative analysis on the instruction difficulty of three datasets—Alpaca, Alpaca-cleaned, and LIMA—employing GPT-4 for scoring. Figures 5, 6 to 7 illustrate their respective score distributions and probability density functions. Alpaca, generated by the Self-Instruct framework, consists of 52K triplets (instruction, input, output). Alpaca-cleaned represents an enhanced version of Alpaca, refined to minimize errors such as incorrect responses and hallucinations. LIMA, by contrast, is composed of 1,000 manually curated instructions. Analysis of score histograms across these figures indicates a predominance of higher scores in LIMA, comparatively fewer in Alpaca, with Alpaca-cleaned exhibiting overall improved scoring relative to Alpaca. These findings corroborate our qualitative assessment of the datasets and affirm the utility of GPT-4 in evaluating instruction difficulty.

### v2

@@ 4.1 Experiment 1: compares the win rates of Phased IFT and One-off IFT
Permutations Self-Instruct WizardLM Koala Vicuna OASST Anthropic Avg. win-rate Base Models
1-2-3 +4.17 +6.42 +5.56 +6.88 +9.31 +14.34 +7.26 Llama2 7B
2-1-3 +1.39 +6.65 +2.30 +6.88 +4.79 +8.91 +4.59
3-1-2 +0.80 +0.94 -2.35 +6.88 +3.83 +8.66 +2.26
1-3-2 +3.57 -1.15 -1.15 +1.88 +2.66 +8.91 +2.14
2-3-1 -2.61 -6.81 -7.65 -5.70 -4.12 -0.39 -4.5846
3-2-1 -7.23 -3.18 -9.09 +1.54 -5.98 +2.73 -4.5847
1-2-3 +1.98 +2.75 +7.78 +11.88 +11.44 +16.28 +7.35 Llama2 13B
2-1-3 +2.33 +1.63 +8.82 +6.39 +6.41 +12.29 +5.57
3-1-2 -0.61 +0.70 +4.39 +5.62 +2.75 +8.40 +2.71
1-3-2 -0.40 +2.14 +6.17 +1.25 +4.72 +9.60 +3.53
2-3-1 -6.83 -10.56 -8.33 -8.12 -3.57 +0.78 -6.44
3-2-1 -5.42 -8.22 -7.69 -8.23 -10.50 -1.97 -7.09
1-2-3 +3.05 +9.67 +9.52 +7.50 +6.32 +12.60 +7.64 Llama3 8B
2-1-3 -2.48 +3.61 +4.94 +3.38 +6.21 +5.24 +3.02
3-1-2 +0.82 +2.12 +0.90 +3.80 +2.53 +6.35 +2.32
1-3-2 -0.41 +0.00 +0.31 -1.97 +5.40 +5.16 +1.40
2-3-1 -11.27 -14.44 -11.70 -14.86 -16.44 -14.22 -13.57
3-2-1 -11.88 -9.39 -5.00 -5.77 -9.23 -3.97 -8.26
Table 5: Ablation studies of various methods on Alpaca-3-stages instruction data. Each entry denotes the win rate of Phased IFT against One-off IFT on the original Alpaca data.
Method Self-Instruct WizardLM Koala Vicuna OASST Anthropic Avg. win-rate Models
Phased IFT stage1 -8.53 -10.32 -15.00 -6.25 -10.11 -4.26 -9.59 Llama2 7B
stage2 +3.57 +2.98 -0.83 +5.00 +3.99 +7.36 +3.34
stage3 +4.17 +6.42 +5.56 +6.88 +9.31 +14.34 +7.26
IFT on stage2 -0.20 +2.52 -0.83 +8.75 +0.53 +0.39 +1.14
IFT on stage3 +0.99 +6.88 +1.67 +16.25 +0.00 +10.08 +4.44
IFT on mixed stage1&2 +2.78 +1.83 -8.33 -5.00 -4.79 +0.00 -1.62
IFT on mixed stage1&3 -2.58 +0.00 -2.50 +3.12 -3.46 -3.88 -1.91
IFT on mixed stage2&3 +0.40 +1.38 +4.17 +8.12 +1.33 +12.79 +3.53
Phased IFT stage1 -6.55 -10.09 -13.61 -6.88 -6.65 -0.78 -7.83 Llama2 13B
stage2 +0.40 -0.23 +4.44 +3.75 +3.72 +16.28 +3.77
stage3 +1.98 +2.75 +7.78 +11.88 +11.44 +16.28 +7.35
IFT on stage2 -2.40 -0.47 +4.09 +3.80 +6.87 +9.06 +2.66
IFT on stage3 +4.56 +7.26 +4.11 +10.12 +3.26 +9.18 +5.80
IFT on mixed stage1&2 -0.60 -1.61 -3.06 +0.62 +2.93 +2.71 -0.09
IFT on mixed stage1&3 +1.39 +0.46 -3.33 +0.00 +6.12 +3.10 +1.33
IFT on mixed stage2&3 +1.39 +3.67 +7.50 +8.12 +7.45 +10.85 +5.68
Experiment 1 demonstrates the effectiveness of Phased IFT compared to One-off IFT, while Experiment 2 reveals that this effectiveness arises from multi-stage sub-datasets with increasing difficulty, not from the multi-stage training itself. Therefore, we designed multi-stage datasets with random segmentation, each segment maintaining the same quantity as the corresponding dataset with increasing difficulty, but the data points were randomly sampled. Table 3 lists the results of three sets of experiments as follows.
The first part of Table 3 presents the win rate of Phased IFT on Alpaca-rand-3-stages relative to One-off IFT on the original Alpaca. Specifically, we used five baseline models, namely Llama2 7/13/70B and Llama3 8/70B. Their performance on six benchmarks was consistently lower than that of One-off IFT, indicated by negative win rates. This suggests that segmenting the dataset randomly into multiple subsets and then performing uptraining did not yield benefits. Similarly, the second part of Table 3 shows the win rate of Phased IFT on Alpaca-cleaned-rand-3-stages relative to One-off IFT on the original Alpaca-cleaned. Based on the results, the win rates of Llama2 7B and Llama3 8B were negative, indicating that multi-stage training did not help.
The third part of Table 3 shows the win rate of Phased IFT on Alpaca-3-stages relative to Phased IFT on Alpaca-rand-3-stages. The results indicate that uptraining on the difficulty-increasing Alpaca-3-stages dataset resulted in significantly higher win rates than uptraining on the randomly segmented Alpaca-rand-3-stages dataset. Specifically, we used five baseline models for uptraining on both Alpaca-3-stages and Alpaca-rand-3-stages. The average win rate of Llama3 70B reached +10.62, Llama2 70B reached +9.71, Llama2 13B reached +8.29, and Llama2 7B and Llama3 8B also reached +5.74 and +5.95, respectively. These models all achieved consistently high win rates, far higher than uptraining on randomly segmented subsets.
These experiments indicate that uptraining on multi-stage datasets with increasing difficulty is beneficial and enhances the model’s capabilities. Uptraining on randomly segmented multi-stage datasets reduced the model’s capabilities, and the results were not as good as single-step fine-tuning on the original dataset (For a detailed graphical presentation, please refer to Appendix A.4). This suggests that multi-stage training methods did not improve the model’s capabilities and yield benefits. The effect of Phased IFT mainly comes from multi-stage subsets with increasing difficulty, which also validates our proposed progressive alignment hypothesis, that aligning pre-trained models to accomplish the final human task is a gradual process.
@@
4.3 Experiment 3: compares the win rates achieved by uptraining on Alpaca-3-stages across all permutations
@@
To investigate the impact of varying the sequence of Alpaca-3-stages on the performance of the Llama-2 7B/13B and Llama3 8B models with clarity and precision, the following structure is proposed. This structure presents the uptraining order across the three stages, denoted as Stage 1 (1), Stage 2 (2), and Stage 3 (3), to illustrate the sequence in which the model was uptrained on different segments of the instruction data:
@@
Table 4 presents the uptraining results of Llama2 7/13B and Llama3 8B models on the three-stages Alpaca-3-stages dataset, evaluating 6 possible permutations. Analyzing the average win rate, a clear pattern emerges: the sequences 1-2-3 and 2-1-3 outperform 1-3-2 and 3-1-2, which in turn outperform 2-3-1 and 3-2-1. This suggests that uptraining on sub-datasets of increasing difficulty can indeed enhance the model’s potential.
The first part of Table 4 details the full permutation experiments on the Llama2 7B model. The 1-2-3 sequence achieved the highest win rates across 6 benchmarks and the average win rate, except for a slight dip below the 2-1-3 sequence in the WizardLM benchmark (e.g., +6.42 vs. +6.65). The average win rate for the 1-2-3 sequence was +7.26, significantly higher than the other permutations. The 2-1-3 sequence also performed well with an average win rate of +4.59, both sequences sharing a common feature of placing the most difficult sub-dataset (stage 3) at the final stage of uptraining. The sequences 3-1-2 and 1-3-2, which place the medium difficulty sub-dataset (stage 2) last, also showed positive win rates of +2.26 and +2.14, respectively. Surprisingly, when the easiest sub-dataset (stage 1) was placed last, the sequences 2-3-1 and 3-2-1 had lower average win rates than One-off IFT on the original Alpaca dataset.
The second part of Table 4 shows the full permutation experiments on the Llama2 13B model. Similarly, the 1-2-3 and 2-1-3 sequences achieved the highest average win rates of +7.35 and +5.57, respectively, significantly outperforming the other permutations. These sequences also shared the characteristic of placing the most difficult sub-dataset (stage 3) last. The 1-3-2 and 3-1-2 sequences achieved average win rates of +3.53 and +2.71, respectively, with the 3-2-1 sequence having the lowest average win rate of -7.09. Interestingly, the permutation experiments on the Llama2 13B model aligned perfectly with the win rate predictions of the progressive alignment hypothesis, ranking the uptraining results from highest to lowest average win rate as follows: 1-2-3 > 2-1-3 > 1-3-2 > 3-1-2 > 2-3-1 > 3-2-1. This further confirms the benefit of uptraining on sub-datasets of increasing difficulty when fine-tuning LLMs.
The third part of Table 4 presents the full permutation experiments on the Llama3 8B model. The 1-2-3 and 2-1-3 sequences achieved the highest average win rates of +7.64 and +3.02, respectively, outperforming the 3-1-2 and 1-3-2 sequences, which had average win rates of +2.32 and +1.40, respectively. Unlike the Llama2 7/13B models, the 2-3-1 sequence had the lowest win rate at -13.57, significantly lower than the -8.26 of the 3-2-1 sequence. Further investigation is needed to understand the reasons behind this result.
In sum, the results indicate a phenomenon where placing more difficult sub-datasets at the final stage of uptraining on differentiated difficulty sub-datasets can significantly enhance the model’s capabilities. This approach markedly increases the model’s win rate, making the generated results closer to human perception, as the judgments of GPT-4 align with human approximately 75%.
4.4 Ablation Studies
We conducted ablation experiments using Llama2 7B/13B on Alpaca-3-stages, incorporating various sub-dataset combinations. These combinations included performing Phased IFT on Alpaca-3-stages, IFT exclusively on stage 2, IFT exclusively on stage 3, IFT on a combined sub-dataset of stage 1 and stage 2, IFT on a combined sub-dataset of stage 1 and stage 3, and IFT on a combined sub-dataset of stage 2 and stage 3. Table 5 presents two sets of ablation experiments, further verifying the role of each sub-dataset in fine-tuning LLMs.
The first part of Table 5 shows the ablation results for the Llama2 7B base model. Performing IFT exclusively on stage 2 yielded a positive gain, with a win rate of +1.14. The win rate for Phased IFT at the stage 2 phase was +3.34, indicating that stage 1 also played a positive role during uptraining. The win rate for IFT exclusively on stage 3 reached +4.44, which was higher than the Phased IFT at stage 2 but lower than the Phased IFT at stage 3. This suggests that stage 3 played a major role in performance improvement during Phased IFT. This result is consistent with the conclusions of Alpagasus Chen et al. (2024), which indicate that high-quality data filtered by GPT is crucial for fine-tuning models. Performing IFT on the combined sub-dataset of stage 1 and stage 2 resulted in a negative win rate, as did performing IFT on the combined sub-dataset of stage 1 and stage 3. This implies that redundant datasets are not conducive to instruction training. However, performing IFT on the combined sub-dataset of stage 2 and stage 3 resulted in a win rate of +3.53, higher than the +1.14 for IFT on stage 2 alone but lower than the +4.44 for IFT on stage 3 alone. This nuanced result suggests that mixing stage 2 data had a detrimental effect, leading to reduced model performance. Nevertheless, when uptraining across all three stages (stage 1, stage 2, and stage 3), the win rate steadily improved, indicating a gradual enhancement in model capability.
This is a very interesting experimental phenomenon, and similar observations were made in ablation experiments with Llama2 13B. This indicates that the model performance improvement achieved through uptraining on increasingly difficult sub-datasets is not a trivial training technique. The quality of the instruction data and the alignment method are key factors in fine-tuning of large language models. This also further supports the hypothesis of alignment progression as a genuine phenomenon.
@@ 5 Conclusion
The paper introduces the progressive alignment hypothesis, and then proposes Phased IFT approach. Utilizing GPT-4 to score the difficulty of instructions, the instruction dataset is categorized into different stages according to the difficulty. Diverging from prior methods that apply a One-off IFT on the whole instruction data, our approach entails staged, supervised uptraining on each stage in ascending order of difficulty. Through extensive experiments, we have validated the progressive alignment hypothesis and the efficacy of Phased IFT, demonstrating its utility in developing diverse instruction dataset and enhancing fine-tuning LLMs.
@@ 6 Broader Impact
This paper discusses a fundamental problem in Large Language Models (LLMs): How to more effectively utilize instruction datasets to align LLMs more efficiently? This paper introduces a simple and effective alignment method, which includes the analysis of instruction difficulty and structuring the dataset into multi-stage sub-datasets, with each sub-dataset increasing in difficulty. Supervised fine-tuning is then carried out through uptraining. We conducted a series of experiments to actively verify the existence of the progressive alignment hypothesis. If the progressive alignment hypothesis is a universal phenomenon, then this work will have a broad impact on the direction of fine-tuning instructions for Large Language Models.
The datasets employed in this work, namely Alpaca 52K and Alpaca-cleaned, are publicly available, and our training framework leverages open-source tools, utilizing Hugging Face’s Trainer and DeepSpeed. Furthermore, we have purchased access to OpenAI’s API to employ GPT-4 for assessing the difficulty of instructions and for evaluating the quality of generated responses. We make all our code and data openly available at https://github.com/xubuvd/PhasedSFT, including training codes, GPT-4 difficulty scoring codes, win-rate calculation codes, and the difficulty-stratified Alpaca and Alpaca-cleaned instruction datasets.
@@ 7 Limitations
Automatic instruction dataset segmentation This study draws heuristic information from the cumulative density curve of difficulty scores in the instruction dataset using human experience, selecting several thresholds to divide the instruction dataset into multiple subsets at different stages, with increasing difficulty levels. Utilizing human experience to select thresholds is a significant task that not only involves partitioning the dataset into subsets from easy to difficult but also requires considering the balance in the quantity of each subset, heavily relying on human expertise. Therefore, a significant challenge lies in developing an algorithm capable of autonomously identifying threshold points from this cumulative density curve, eliminating the dependence on human intuition, and ensuring a balanced quantity across the multi-stage subsets
@@ Appendix A Appendix
Table 6: Comparison of percentage distribution of difficulty scores on the Alpaca52K dataset with GPT-3.5 and GPT-4
GPT [1.0, 2.5) [2.5, 3.5] (3.5, 4.5) [4.5, 5.0]
3.5 2.18% 62.28% 25.96% 9.56%
4 76.3% 13.38% 4.1% 6.17%
@@
Figure 8: Comparison of the win rate trends for uptraining on multi-stage sub-datasets with increasing difficulty (solid line) and uptraining on randomly divided multi-stage sub-datasets (dashed line). The gray horizontal line in the middle represents the baseline of One-off IFT on the original dataset.
A.1 Comparison of percentage distribution of difficulty scores on the Alpaca52K dataset using GPT-3.5 and GPT-4
In our initial assessment, we compared the differences of GPT-3.5 and GPT-4 (gpt-4-0613) in assigning difficulty scores to the Alpaca 52K dataset through human evaluation, the score ranges from 1.0 to 5.0. We divided the scores between 1.0 and 5.0 into four intervals, with the quantity percentage of each interval shown in the Table 6.
The distribution of scores reveal a noticeable regression toward the mean for GPT-3.5, with scores between 2.5 and 3.5 accounting for 62.28% of the total. A manual review of 200 samples indicated a mismatch between the assigned difficulty levels and human judgment, particularly noting an excess of simpler questions within the dataset. In contrast, GPT-4’s difficulty ratings aligned with human evaluations for approximately 75% of the cases, with scores ranging between 2.5 and 3.5 comprising 13.38% of the total.
A.2 Histogram and density curve for Alpaca, Alpaca-cleaned and LIMA
We conducted a comparative analysis on the instruction difficulty of three datasets—Alpaca, Alpaca-cleaned, and LIMA—employing GPT-4 for scoring. Figures 4, 5 to 6 illustrate their respective score distributions and probability density functions. Alpaca, generated by the Self-Instruct framework, consists of 52K triplets (instruction, input, output). Alpaca-cleaned represents an enhanced version of Alpaca, refined to minimize errors such as incorrect responses and hallucinations. LIMA, by contrast, is composed of 1,000 manually curated instructions. Analysis of score histograms across these figures indicates a predominance of higher scores in LIMA, comparatively fewer in Alpaca, with Alpaca-cleaned exhibiting overall improved scoring relative to Alpaca. These findings corroborate our qualitative assessment of the datasets and affirm the utility of GPT-4 in evaluating instruction difficulty.